OpenStudy (nuccioreggie):

The number line shows the distance in meters of two birds, A and B, from a worm located at point X: A horizontal number line extends from negative 3 to positive 3. The point labeled as A is at negative 2.5, the point 0 is labeled as X, and the point labeled B is at 2.5. Write an expression using subtraction to find the distance between the two birds. (5 points) Show your work and solve for the distance using additive inverses

OpenStudy (nuccioreggie):

@amistre64

OpenStudy (nuccioreggie):

@dan815

OpenStudy (nuccioreggie):

@GirlgoyleH.

OpenStudy (girlgoyleh.):

I'm no good at math, sorry

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

we have the subsequent drawing: 

OpenStudy (michele_laino):

the requested distance d, is given by the subsequent formula: \[\Large d = B - A = ...?\]

OpenStudy (nuccioreggie):

Ok so i have to write the answer in the answe box bow do i write it because i cant write a numberline???

OpenStudy (michele_laino):

hint: \[\Large d = B - A = 2.5 - \left( { - 2.5} \right) = ...?\]

OpenStudy (nuccioreggie):

0??

OpenStudy (michele_laino):

no, it is: \[\Large 2.5 - \left( { - 2.5} \right) = 2.5 + 2.5 = ...?\]

OpenStudy (michele_laino):

since: - * - = +

OpenStudy (nuccioreggie):

5

OpenStudy (michele_laino):

that's right!

OpenStudy (nuccioreggie):

Ok so can you write out for me how im going to write the answer in the answer box and thank you

OpenStudy (michele_laino):

you have to write: d=5

OpenStudy (nuccioreggie):

Write it in the answer box

OpenStudy (michele_laino):

5

OpenStudy (nuccioreggie):

Right d=5 ? And how did you get it

OpenStudy (michele_laino):

since I applied this formula: d=B-A where A= -2.5, and B=2.5

OpenStudy (nuccioreggie):

So thats going to be what i right and thats all?

OpenStudy (michele_laino):

yes! we can solve your exercise, using fraction, like below: \[d = B - A = \frac{5}{2} - \left( { - \frac{5}{2}} \right) = \frac{5}{2} + \frac{5}{2} = \frac{{5 + 5}}{2} = \frac{{10}}{2} = 5\]

OpenStudy (michele_laino):

fractions*

OpenStudy (nuccioreggie):

Ok thanks. Can YOU. Help me with 3 more please?

OpenStudy (michele_laino):

ok!